OpenStudy (anonymous):

Find the point on the terminal side of θ = negative three pi divided by four that has an x coordinate of -1. Show your work for full credit.

zepdrix (zepdrix):

Hey there!

zepdrix (zepdrix):

\[\Large\rm \theta=-\frac{3\pi}{4},\qquad x=-1\]Hmm, let's see...

zepdrix (zepdrix):